OpenStudy (anonymous):

The ratio of similarity of the following hexagons is 3/4. What is a/b?

OpenStudy (anonymous):

the side length in the first hexagon is 3a+b and the second a+2b

OpenStudy (anonymous):

answers: 3/7, 1/7, 2/9, 1/9

OpenStudy (ash2326):

We are given the ratio of similarity as \[\frac{3}{4}\] Side length of first hexagon is 3a+b and for the second is a+2b so \[\frac{3a+b}{a+2b}=\frac{3}{4}\] Cross multiply, we get \[4\times (3a+b)=3 \times (a+2b)\] \[12a+4b=3a+6b\] Subtract both sides by 3a and 4b \[9a=2b\] divide both sides by 9b, we get \[\frac{a}{b}=\frac{2}{9}\]

OpenStudy (anonymous):

oh that makes sense...thank you!

OpenStudy (ash2326):

Welcome:D